OpenStudy (anonymous):

solve the quadratic equation ?? x2 – 6x – 16 = 0

OpenStudy (lgbasallote):

use the quadratic formula \[x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\] do you know how to use it?

OpenStudy (anonymous):

yournot suppose to use that!!

OpenStudy (anonymous):

\[x ^{2}-8x+2x-16 = 0\] \[x(x-8)+2(x-8) =0\] (x+2)(x-8)=0

OpenStudy (zzr0ck3r):

x^2-6x-16=0 (x-3)^2 = 25 x-3=+-sqrt(25) x-3 = +-5 x = -5+3 or x = 5+3 x = -2 or x =8 now you see all three methods for solving this:)

OpenStudy (anonymous):

yes thanks